OpenStudy (anonymous):

how to diferenciate de equation y=ln(1+x^x)^e^x ?

OpenStudy (anonymous):

I believe the problem here is how do I deal with the x^x term. Try use x^x = exp(xlnx) and apply chain rule, this may help.

OpenStudy (anonymous):

1 -------------- . (eˆx).[ (1+xˆx)ˆ(eˆx-1) ] . x.xˆ(x-1) (1+xˆx)ˆ(eˆx)

OpenStudy (anonymous):

\[ ( 1\div (1+x ^{x})^{eˆx}) \times (e ^{x}.(1+x ^{x})^{e ^{x-1}})\times (x.x ^{x-1})\]

OpenStudy (anonymous):

thx for support!